starting from rest, a train is uniformly accelerated and reaches a velocity of 20km/h after 15cesconds. what is its acceleration


1
Expert's answer
2022-04-06T14:02:50-0400

Answer

Acceleration is

"a=\frac{v}{t}\\=\frac{20000}{3600*15}\\=0.37m/s^2"